a student eats dinner rated at 2000 kilo calories . he wishes to do equivalent amount of work in the gynasium by lifting a 50 kg mass. how many times must he raise the weight to expand this much energy ? assume that he raise s the weight a distance of 2 m each time and that no work is done when the weight is dropped at the floor....
1
Expert's answer
2012-10-03T12:08:06-0400
The work done to raise 50 kg mass a distance of 2 m is

A = MgH = 50[kg]*9.8[m/s²]*2[m] = 980[J].

1[c] = 4.184[J], so 2000*10³[c] = 8368*10³[J]. Therefore, a student need to do

N = 8368*10³[J]/980[J] ≈ 8538.77 raisings.